OpenStudy (megannicole51):

Find the first 4 terms of the taylor series for the function about the point a. cos(theta) a=pi/4

OpenStudy (megannicole51):

i know cos(x)=1-x^2/2!+x^4/4!-x^6/6!+x^8/8!

OpenStudy (megannicole51):

can i just plug in a?

OpenStudy (megannicole51):

@dumbcow

OpenStudy (dumbcow):

not quite, that series assumes "a=0"

OpenStudy (megannicole51):

how do u know..then what do i do with a=pi/4

OpenStudy (dumbcow):

\[= \cos(a)-\sin(a)(x-a)-\frac{\cos(a)}{2}(x-a)^{2}+\frac{\sin(a)}{6}(x-a)^{3}\] now just plug in "a = pi/4"

OpenStudy (megannicole51):

you are my savior tonight i hope you know!

OpenStudy (dumbcow):

well thank you :) ...but you can do it without me just use the taylor series formula and do the derivatives

OpenStudy (megannicole51):

im starting to get the hang of it...kinda lol